# Customized Interior-Point Methods Solver for Embedded Real-Time Convex Optimization

### v1

@@ Abstract
This paper presents a customized convex optimization solver tailored for embedded real-time optimization, which frequently arise in modern guidance and control (G&C) applications. The solver employs a practically efficient predictor-corrector type primal-dual interior-point method (PDIPM) combined with a homogeneous embedding framework for infeasibility detection. Unlike conventional homogeneous self-dual embedding formulations, the adopted approach can directly handle quadratic cost functions without requiring problem reformulation. To support a systematic workflow, we also develop a code generation tool that analyzes the sparsity pattern of the provided problem family and generates customized solver code using a predefined code template. The generated solver code is written in C with no external dependencies other than the standard library math.h, and it supports complete static allocation of all data. Additionally, it provides parsing information to facilitate the use of the solver API by end users. Benchmark results and numerical experiments on an embedded platform demonstrate that the developed solver outperforms existing solvers in both efficiency and reliability.
Index Terms:
Convex optimization, Customized solver, Embedded real-time optimization, Interior-point methods
@@ I Introduction
Optimization has become an important tool in modern guidance and control (G&C) applications [1]. By transcribing the G&C problem into a constrained optimization problem, one can efficiently obtain desired solutions while accounting for complex dynamics, actuator limitations, and operational constraints. Among various approaches, convex optimization has emerged as a prominent approach due to its favorable theoretical and practical properties for onboard real-time applications, such as guaranteed convergence to a global optimum and existence of polynomial-time algorithms [2, 3]. As a result, it plays a key role in enabling practical implementation of optimization-based techniques, such as model predictive control (MPC) [4, 5, 6, 7] and onboard trajectory optimization [8, 9, 10]. It also frequently arises as subproblems in sequential methods for nonconvex problems, such as sequential quadratic programming (SQP) [11] and sequential convex programming (SCP) [12, 13].
@@
However, the computational bottleneck associated with solving optimization problems onboard remains a major challenge to the practical implementation of these methods. Therefore, an efficient and reliable solver is essential for the successful deployment. Several numerical solution techniques have been developed over the past decades. In particular, interior-point methods (IPMs) [14, 15] and alternating direction method of multipliers (ADMM) [16] are among the most popular choices. Both algorithms can effectively handle general convex conic programs, including linear programs (LPs), quadratic programs (QPs), and second-order cone programs (SOCPs), which encompass most G&C problems [17]. ADMM is particularly attractive for large-scale problems due to its low computational cost per iteration. However, it is known to suffer from poor tail convergence at high precision [18, 16]. It can also be sensitive to the ill-conditioned problem data, potentially requiring careful tuning of the parameters, such as step size, prior to execution [19]. These limitations may hinder reliable real-time embedded operation, and may adversely affect the convergence of sequential methods such as SQP and SCP [20].
@@
On the other hand, IPMs are capable of achieving high-precision solutions in far fewer iterations, almost independent of problem size or conditioning. Moreover, since most G&C problems are typically formulated as small- to medium-scale problems, IPMs tend to exhibit strong performance in such applications [21, §1]. They have been implemented in several modern convex optimization solvers, including CVXGEN [22], QOCO [23], ECOS [24], MOSEK [25], and Clarabel [26]. However, since CVXGEN and QOCO rely solely on standard IPM implementations, which assume the existence of a strictly feasible solution, their algorithms fail to converge when solving infeasible problem instances. Such infeasibility scenarios are commonly encountered in practical G&C applications. For example, instances of the powered descent guidance problem [27, 28] may become infeasible depending on the specified time-of-flight. To ensure reliable G&C operation, it is important to detect infeasibility in problem instances within a finite number of iterations. A widely adopted approach to address this challenge is homogeneous embedding [29, 30, 31], which reformulates the problem into a slightly larger feasibility problem. This lifted problem is always feasible, which ensures that the IPM algorithm converges within a finite number of iterations regardless of the feasibility of the original problem. Once a solution is found from the lifted problem, it can be used either to recover the optimal solution to the original problem or to provide a certificate of infeasibility, depending on the status of the solution. In the case of ECOS and MOSEK, a homogeneous self-dual embedding [29, 32] formulation is employed to enable infeasibility detection. However, this formulation is limited to problems with linear cost functions due to the reliance on the self-dual formulation. Thus, to apply such a homogeneous embedding to problems with quadratic cost functions, the original problem needs to be reformulated as a linear objective problem by introducing a slack variable and an additional second-order cone constraint [26]. This transformation leads to unnecessary increases in problem size and degrades the sparsity of the problem data, which can adversely affect solver performance. To address this issue, previous studies [30, 33, 31] have proposed a homogeneous embedding formulation capable of handling quadratic cost functions. Clarabel is one of the few solvers that adopt this technique within an IPM framework and has demonstrated its effectiveness across diverse problem classes. Furthermore, it supports more general conic constraints beyond SOCPs, including semidefinite cones, exponential cones and power cones. However, it is designed as a general-purpose solver, rather than tailored for embedded applications. In particular, its implementation in Julia and Rust poses practical challenges for deployment on embedded systems. Although Rust is capable of being used in embedded environments and ensures memory safety through its ownership model, its ecosystem is still far less mature than that of C. As a result, limitations exist in terms of compatibility with existing C-based legacy systems and the availability of library support.
@@
Another important factor in embedded real-time optimization is effectively exploiting the sparsity of problem data. Embedded real-time applications typically involve repeatedly solving problems with the same structure and sparsity pattern but with different data instances. This feature makes a custom solver particularly well-suited for such applications. A custom solver refers to a solver specifically tailored to the structure of the target problem class. It exploits the known structure and sparsity pattern to precompute certain information, such as the sparsity information for matrix factorization, during an offline phase. The precomputed information is then utilized during online operation, thereby improving computational efficiency. Representative custom solvers such as CVXGEN and QOCO employ custom linear algebra routines in which matrix operations and factorizations are explicitly hard-coded. While this approach enables faster operation by reducing overhead during execution, it has the drawback that the resulting code size increases rapidly with the problem size, limiting its scalability beyond small-scale problems [34]. This issue can be mitigated by implementing the numerical routines for linear algebra using a loop-based structure, while caching only the sparsity information obtained offline [35, 36]. A further distinctive feature of a customized solver is its ability to statically allocate all data. In the case of general-purpose solvers, dynamic memory allocation is inevitable in order to accommodate arbitrary problem structures. However, dynamic memory allocation introduces risks such as memory leaks and fragmentation, which are particularly undesirable in embedded systems where computational and memory resources are typically limited. Furthermore, it complicates the verification and validation process due to uncertainties introduced during the allocation process. In contrast, a customized solver can statically allocate all problem data by leveraging the known problem structure in advance, making it well-suited for embedded applications.
@@
  • The solver developed in this study employs an efficient and robust algorithmic formulation that combines the predictor-corrector type PDIPM with a homogeneous embedding model capable of directly handling quadratic objective terms. This algorithmic structure allows efficient treatment of a wide range of problem classes without reformulation, thereby avoiding unnecessary performance degradation.
  • A solver customization framework is applied to enable efficient execution in embedded system environments. Moreover, a dedicated code generation tool is developed to reduce user workload by automating the customization process. The generated code has no external dependencies other than the standard library math.h, and its size remains nearly constant regardless of problem dimensions due to the use of loop-based structures for linear algebra routines. It also supports complete static allocation of problem data, and is designed for memory efficiency by utilizing only the upper triangular part of a symmetric matrix. In addition, the code generation tool provides parsing information to facilitate the use of the solver API, significantly reducing the user workload for handling problem data in C-based environments.
  • Extensive performance evaluations are conducted to demonstrate the effectiveness of the developed solver. In benchmark tests, its performance and robustness are compared against existing solvers across various problem sets. Furthermore, numerical experiments on an embedded system validate its capability for real-time optimization in resource-constrained environments.
@@ II Second-Order Cone Programming with a Quadratic Cost Function
+:=$\mathbb{R}_{+}:=\left\{x\in\mathbb{R}\left.\right|\ x\geq 0\right\}$
the second-order cone (SOC) for $d>1$,
$\mathbb{Q}^{d}:=\left\{(x_{0},x_{1})\in\mathbb{R}\times\mathbb{R}^{d-1}\left.\right|\ \|x_{1}\|_{2}\leq x_{0}\right\}$
or both. Without loss of generality, it is further assumed that the cone $\mathcal{K}$ is ordered such that the first $l\geq 0$ cones in $\mathcal{K}$ correspond to the NNC, followed by $n_{SOC}\geq 0$ SOCs, which is given by
$\mathcal{K}=\mathbb{R}_{+}^{(1)}\times\cdots\times\mathbb{R}_{+}^{(l)}\times\mathbb{Q}^{d_{1}}\times\cdots\times\mathbb{Q}^{d_{n_{SOC}}}$ (1)
The optimal solution to the primal problem $\mathcal{P}$, if it exists, is denoted as $(x^{\ast},s^{\ast})$, and the corresponding optimal value as $p^{\ast}$.
@@
$\mathcal{K}^{\ast}=\{y\left.\right|\ \langle y,x\rangle\geq 0,\ \forall x\in\mathcal{K}\}$
The optimal solution to the dual problem $\mathcal{D}$, if it exists, is denoted as $(y^{\ast},z^{\ast})$, with the corresponding optimal value $d^{\ast}$.
@@
Therefore, finding a solution pair $(x,s,y,z)$ that satisfies the KKT conditions in (2) corresponds to obtaining the optimal primal-dual solution pair $(x^{\ast},s^{\ast},y^{\ast},z^{\ast})$.
@@ III-A Interior-Point Methods
$\det{(x)}=\begin{cases}x&x\in\mathbb{R}_{+}\\ x_{0}^{2}-x_{1}^{T}x_{1}&x\in\mathbb{Q}^{d}\end{cases}$
Here, $N=l+n_{SOC}$ denotes the order of the cone $\mathcal{K}$.
@@
The corresponding optimality conditions for the problem (4) are given by [38]:
@@ III-B Homogeneous Embedding for Quadratic Cost Functions
Note that the quadratic objective term is explicitly accounted for in the formulation. As a result, unlike the conventional homogeneous self-dual embedding [29, 41], the given HMCP does not require problem reformulation with additional slack variables and SOC constraints as in (7). This prevents unnecessary degradation in solver performance caused by increased problem size and reduced sparsity. Moreover, since the HMCP becomes identical to the homogeneous self-dual embedding when $Q=0$, it can be regarded as a more general homogeneous embedding formulation.
$\begin{split}\underset{x,s}{\text{minimize}}\quad&q^{T}x+t\\ \text{subject to}\quad&Ax=b,\quad Gx+s=h\\ &\left\|\begin{matrix}t-\frac{1}{2}\\ Q^{1/2}x\end{matrix}\right\|_{2}\leq t+\frac{1}{2}\\ &s\in\mathcal{K},\quad t\geq 0\end{split}$ (7)
@@ III-C Nesterov-Todd Scaling
    SOC : For $s,\ z\in\mathbb{Q}^{d}$ with $d>1$, let $\bar{s}=(\bar{s}_{0},\bar{s}_{1})$, $\bar{z}=(\bar{z}_{0},\bar{z}_{1})\in\mathbb{R}\times\mathbb{R}^{d-1}$ be the normalized points
@@
    and $\gamma$, $\bar{w}$, $\eta$ be
    $\gamma=\left(\frac{1+\bar{s}^{T}\bar{z}}{2}\right)^{1/2}$
    $\bar{w}=\frac{1}{2\gamma}\left(\bar{s}+\begin{bmatrix}\bar{z_{0}}\\ -\bar{z}_{1}\end{bmatrix}\right)$
    $\eta=\left(\frac{s_{0}^{2}-s_{1}^{T}s_{1}}{z_{0}^{2}-z_{1}^{T}z_{1}}\right)^{1/4}$
@@
$W=\texttt{blkdiag}(W_{1},\cdots,W_{D})$
@@ III-D Search Directions
The solver developed in this paper utilizes PDIPMs, which are known to be among the most efficient frameworks within IPMs [45]. PDIPMs compute a Newton-like search direction by linearizing the (scaled) central path equations and loosely follow the central path toward the solution. Since the iterates of PDIPMs do not lie exactly on the central path (i.e., equations (6b), (6c) or (10) are not precisely satisfied for the barrier parameter $\beta$), PDIPMs introduce the duality measure $\mu$ as an indicator of the approximated duality gap at each point in the search space instead of $\beta$ [45].
$\mu=\frac{s^{T}z+\kappa\tau}{D+1}$
@@
$(W^{-T}s)\circ(Wz)=\sigma\mu e,\quad\kappa\tau=\sigma\mu e$
@@
$\Lambda:=\begin{bmatrix}Q&A^{T}&G^{T}&q\\ -A&0&0&b\\ -G&0&0&h\\ -q^{T}-\frac{2}{\tau}x^{T}Q&-b^{T}&-h^{T}&\frac{1}{\tau^{2}}x^{T}Qx\end{bmatrix}$
@@
$\mu_{a}=\frac{(s+\alpha_{a}\Delta s_{a})^{T}(z+\alpha_{a}\Delta z_{a})+(\kappa+\alpha_{a}\Delta\kappa_{a})(\tau+\alpha_{a}\Delta\tau_{a})}{D+1}$
and $\alpha_{a}$ is the step size determined via line search given in (17) with $\Delta(s,z,\kappa,\tau)=\Delta(s_{a},z_{a},\kappa_{a},\tau_{a})$.
$\begin{split}\alpha=\sup\{\alpha\in(0,1)\left.\right|\ &(s,z)+\alpha(\Delta s,\Delta z)\in\mathcal{K}\times\mathcal{K}^{\ast},\\ &(\kappa,\tau)+\alpha(\Delta\kappa,\Delta\tau)\geq 0\}\end{split}$ (17)
@@
where the step size $\alpha$ is determined via line search as given in (17) using $\Delta(s,z,\kappa,\tau)=\Delta(s_{c},z_{c},\kappa_{c},\tau_{c})$.
@@
$\begin{split}\Delta\xi&:=\begin{bmatrix}\Delta x^{T}&\Delta y^{T}&\Delta z^{T}\end{bmatrix}^{T}\\ b_{1}&:=\begin{bmatrix}-q^{T}&b^{T}&h^{T}\end{bmatrix}^{T}\\ b_{2}&:=\begin{bmatrix}-d_{x}^{T}&d_{y}^{T}&(d_{z}-W(\lambda\backslash d_{s}))^{T}\end{bmatrix}^{T}\end{split}$
@@
$K:=\begin{bmatrix}Q&A^{T}&G^{T}\\ A&0&0\\ G&0&-W^{T}W\end{bmatrix}$
@@
$\Delta\xi=\Delta\xi_{2}+\Delta\tau\Delta\xi_{1}$
@@
$\Delta s$ $=W(\lambda\backslash d_{s}-W\Delta z)$
$\Delta\kappa$ $=(d_{\kappa}-\kappa\Delta\tau)/\tau$
@@
To further improve computational efficiency in matrix factorization, we adopt the structure exploitation strategy for SOC scaling proposed in [24] to obtain a sparse KKT matrix. For a SOC $\mathbb{Q}^{d}$, the scaling matrix $W_{SOC}$ in (12b) propagates to the (3, 3) block in the KKT matrix as $-W^{T}_{SOC}W_{SOC}$, which is a dense matrix with $d^{2}$ nonzero entries. This results in a significant number of nonzeros in both the KKT matrix and its factorization, even for problems with sparse data. By exploiting the diagonal-plus-rank-one structure of $W_{SOC}$, we can expand $W^{T}_{SOC}W_{SOC}$ to a sparse form as
$\tilde{W}_{SOC}^{T}\tilde{W}_{SOC}=\eta^{2}\begin{bmatrix}\bar{D}&v&u\\ v^{T}&1&0\\ u^{T}&0&-1\end{bmatrix}$
@@
$\bar{D}:=\begin{bmatrix}a&\\ &I_{n-1}\end{bmatrix},\ u:=\begin{bmatrix}u_{0}\\ u_{1}\bar{w}_{1}\end{bmatrix},\ v:=\begin{bmatrix}0\\ v_{1}\bar{w}_{1}\end{bmatrix}$
@@
Moreover, the expanded $\tilde{W}^{T}_{SOC}\tilde{W}_{SOC}$ is quasi-definite [46, 24], and consequently, the (regularized) KKT matrix also becomes quasi-definite. This property ensures a stable $LDL^{T}$ factorization of the KKT matrix for any permutation. Details on regularization and $LDL^{T}$ factorization of the KKT matrix are provided in section III-E.

### v2

@@ Abstract
This paper presents a customized second-order cone programming (SOCP) solver tailored for embedded real-time optimization, which frequently arises in modern guidance and control (G&C) applications. The solver employs a practically efficient predictor-corrector type primal-dual interior-point method (PDIPM) combined with a homogeneous embedding framework for infeasibility detection. Unlike conventional homogeneous self-dual embedding formulations, the adopted approach can directly handle quadratic cost functions without requiring problem reformulation. This capability allows the solver to directly address quadratic objective SOCP problems, while avoiding unnecessary performance degradation caused by the loss of sparsity due to problem reformulation. To support a systematic workflow, we also develop a code generation tool that analyzes the sparsity pattern of the problem to be solved and generates customized solver code using a predefined code template. The generated solver code is written in C with no external dependencies other than the standard library math.h, and it supports complete static allocation of all data. Additionally, it provides parsing information to facilitate the use of the solver by end users. Finally, benchmark and numerical experiments on an embedded platform demonstrate that the developed solver outperforms the existing solvers on problem scales typical of G&C applications.
@@ I Introduction
Optimization has become an important tool in modern guidance and control (G&C) applications [26]. By transcribing the G&C problem into a constrained optimization problem, one can efficiently obtain desired solutions while accounting for complex dynamics, actuator limitations, and operational constraints. Among various approaches, convex optimization has emerged as a prominent approach due to its favorable theoretical and practical properties for onboard real-time applications, such as guaranteed convergence to a global optimum and existence of polynomial-time algorithms [9, 59]. As a result, it plays a key role in enabling practical implementation of optimization-based techniques, such as model predictive control (MPC) [56, 17, 18, 30] and onboard trajectory optimization [35, 1, 49, 6, 14, 57, 33, 58]. It also frequently arises as subproblems in sequential methods for nonconvex problems, such as sequential quadratic programming (SQP) [7] and sequential convex programming (SCP) [37, 8].
@@
However, solving optimization problems onboard remains a major challenge to the practical implementation of these methods due to the stringent real-time execution requirements under embedded hardware with limited computational and memory resources [26, 22, 14]. Therefore, an efficient and reliable solver is essential for the successful deployment. Several numerical solution techniques have been developed over the past decades. In particular, interior-point methods (IPMs) [45, 60] and alternating direction method of multipliers (ADMM) [10] are among the most popular choices. Both algorithms can effectively handle general convex conic programs, including linear programs (LPs), quadratic programs (QPs), and second-order cone programs (SOCPs), which encompass most G&C problems [36]. ADMM is particularly attractive for large-scale problems due to its low computational cost per iteration. However, it is known to suffer from poor tail convergence at high precision [29, 10]. It can also be sensitive to the ill-conditioned problem data, potentially requiring careful tuning of the parameters, such as step size, prior to execution [52]. These limitations may hinder reliable real-time embedded operation, and may adversely affect the convergence of sequential methods such as SQP and SCP [32].
@@
On the other hand, IPMs are capable of achieving high-precision solutions in far fewer iterations, almost independent of problem size or conditioning. Moreover, since most G&C problems are typically formulated as small- to medium-scale problems, IPMs tend to exhibit strong performance in such applications [48, §1]. They have been implemented in several modern convex optimization solvers, including CVXGEN [39], QOCO [13], ECOS [20], MOSEK [4], and Clarabel [27]. However, since CVXGEN and QOCO rely solely on standard IPM implementations, which assume the existence of a strictly feasible solution, their algorithms fail to converge when solving infeasible problem instances. Such infeasibility scenarios are commonly encountered in practical G&C applications. For example, instances of the powered descent guidance problem may become infeasible depending on the specified time-of-flight [1, 22]. To ensure reliable G&C operation, it is important to detect infeasibility in problem instances within a finite number of iterations. A widely adopted approach to address this challenge is homogeneous embedding [62, 3, 63], which reformulates the problem into a slightly larger feasibility problem. This lifted problem is always feasible, which ensures that the IPM algorithm converges within a finite number of iterations regardless of the feasibility of the original problem. Once a solution is found from the lifted problem, it can be used either to recover the optimal solution to the original problem or to provide a certificate of infeasibility, depending on the status of the solution. In the case of ECOS and MOSEK, a homogeneous self-dual embedding [62, 54] formulation is employed to enable infeasibility detection. However, this formulation is limited to problems with linear cost functions due to the reliance on the self-dual formulation. Thus, to apply such a homogeneous embedding to problems with quadratic cost functions, the original problem needs to be reformulated as a linear objective problem by introducing a slack variable and an additional second-order cone constraint [27]. This transformation leads to unnecessary increases in problem size and degrades the sparsity of the problem data, which can adversely affect solver performance. To address this issue, previous studies [3, 47, 63] have proposed a homogeneous embedding formulation capable of handling quadratic cost functions. Clarabel is one of the few solvers that adopt this technique within an IPM framework and has demonstrated its effectiveness across diverse problem classes. Furthermore, it supports more general conic constraints beyond SOCPs, including semidefinite cones, exponential cones and power cones. However, it is designed as a general-purpose solver, rather than tailored for embedded applications. In particular, its implementation in Julia and Rust poses practical challenges for deployment on embedded systems. Although Rust is capable of being used in embedded environments and ensures memory safety through its ownership model, its ecosystem is still far less mature than that of C. As a result, limitations exist in terms of compatibility with existing C-based legacy systems and the availability of library support.
@@
Another important factor in embedded real-time optimization is effectively exploiting the sparsity of problem data. Embedded real-time applications typically involve repeatedly solving problems with the same structure and sparsity pattern but with different data instances. This feature makes a custom solver particularly well-suited for such applications. A custom solver refers to a solver specifically tailored to a set of problems that share a common structure and sparsity pattern. It exploits the known structure and sparsity pattern to precompute certain information, such as the sparsity information for matrix factorization, during an offline phase. The precomputed information is then utilized during online operation, thereby improving computational efficiency. Representative customized IPM solvers such as CVXGEN and QOCO employ custom linear algebra routines in which matrix operations and factorizations are explicitly hard-coded. While this approach enables faster operation by reducing overhead during execution, it has the drawback that the resulting code size increases rapidly with the problem size, limiting its scalability beyond small-scale problems [23]. This issue can be mitigated by implementing the numerical routines for linear algebra using a loop-based structure, while caching only the sparsity information obtained offline [21, 5]. A further distinctive feature of a customized solver is its ability to statically allocate all data. In the case of general-purpose solvers, dynamic memory allocation is inevitable in order to accommodate arbitrary problem structures. However, dynamic memory allocation introduces risks such as memory leaks and fragmentation, which are particularly undesirable in embedded systems where computational and memory resources are typically limited. Furthermore, it complicates the verification and validation process due to uncertainties introduced during the allocation process. In contrast, a customized solver can statically allocate all problem data by leveraging the known problem structure in advance, making it well-suited for embedded applications.
@@
  • The solver developed in this study employs a PDIPM capable of infeasibility detection for a broad class of SOCPs, including LPs, QPs, SOCPs with linear objectives and even those with quadratic objectives. This algorithmic formulation allows efficient treatment of a wide range of problems without reformulation, thereby avoiding unnecessary performance degradation.
  • A solver customization framework is applied to enable efficient execution in embedded system environments. Moreover, a dedicated code generation tool is developed to reduce user workload by automating the customization process. The generated code has no external dependencies other than the standard library math.h, and its size remains nearly constant regardless of problem dimensions due to the use of loop-based structures for linear algebra routines. It also supports complete static allocation of problem data, and is designed for memory efficiency by utilizing only the upper triangular part of a symmetric matrix. In addition, the code generation tool provides parsing information to facilitate the use of the solver, significantly reducing the user workload for handling problem data in C-based environments.
  • Extensive performance evaluations are conducted to demonstrate the effectiveness of the developed solver. Its performance and robustness are compared against existing solvers across standard benchmark problem sets. Furthermore, numerical experiments on an embedded system using G&C examples validate its capability for real-time optimization in resource-constrained environments.
@@ II Second-Order Cone Programming with a Quadratic Cost Function
$\mathbb{R}_{+}:=\left\{x\in\mathbb{R}\left.\right|\ x\geq 0\right\}$ (1)
the second-order cone (SOC) with dimension $d>1$,
$\mathbb{Q}^{d}:=\left\{(x_{0},x_{1})\in\mathbb{R}\times\mathbb{R}^{d-1}\left.\right|\ \|x_{1}\|_{2}\leq x_{0}\right\}$ (2)
or both. Without loss of generality, it is further assumed that the cone $\mathcal{K}$ is ordered such that the first $l\geq 0$ cones in $\mathcal{K}$ correspond to the NNC, followed by $n_{SOC}\geq 0$ SOCs with dimension $d_{i}$ $(i=1,2,\cdots,n_{SOC})$, which is given by
$\mathcal{K}=\mathbb{R}_{+,1}\times\cdots\times\mathbb{R}_{+,l}\times\mathbb{Q}^{d_{1}}\times\cdots\times\mathbb{Q}^{d_{n_{SOC}}}$ (3)
In the case where $l=0$ or $n_{SOC}=0$, the corresponding constituent cones (i.e., the NNC blocks or the SOC blocks) are omitted in the representation of $\mathcal{K}$ in (3). The optimal solution to the primal problem $\mathcal{P}$, if it exists, is denoted as $(x^{\ast},s^{\ast})$, and the corresponding finite optimal value as $p^{\ast}$.
@@
$\mathcal{K}^{\ast}=\{y\left.\right|\ \langle y,x\rangle\geq 0,\ \forall x\in\mathcal{K}\}$ (4)
The optimal solution to the dual problem $\mathcal{D}$, if it exists, is denoted as $(y^{\ast},z^{\ast})$, with the corresponding finite optimal value $d^{\ast}$.
@@
Therefore, obtaining the primal-dual optimum $(x^{\ast},s^{\ast},y^{\ast},z^{\ast})$ reduces to finding a primal-dual feasible $(x,s,y,z)$ with zero duality gap, since $s^{T}z$ in (5d) is identical to the duality gap for primal-dual feasible $(x,s,y,z)$ [9, §5.5.3].
@@ III-A Interior-Point Methods
$\det{(x)}=\begin{cases}x&x\in\mathbb{R}_{+}\\ x_{0}^{2}-x_{1}^{T}x_{1}&x\in\mathbb{Q}^{d}\end{cases}$ (7)
@@
where the prefix $\text{int}(\cdot)$ represents the interior of the corresponding cone. The corresponding optimality conditions for the problem (9) are given by [25]:
@@ III-B Homogeneous Embedding for Quadratic Cost Functions
Note that the quadratic objective term is explicitly accounted for in the formulation. As a result, unlike the conventional homogeneous self-dual embedding [62, 42], the presented HMCP does not require problem reformulation that introduces an additional slack variable $r$ and a SOC constraint to replace the quadratic objective term, as in (12) [34]. This prevents unnecessary degradation in solver performance caused by increased problem size and reduced sparsity. Moreover, since the HMCP becomes identical to the homogeneous self-dual embedding when $Q=0$, it can be regarded as a more general homogeneous embedding formulation.
$\begin{split}\underset{x,s,r}{\text{minimize}}\quad&q^{T}x+r\\ \text{subject to}\quad&Ax=b,\quad Gx+s=h\\ &\left\|\begin{matrix}r-\frac{1}{2}\\ Q^{1/2}x\end{matrix}\right\|_{2}\leq r+\frac{1}{2}\\ &s\in\mathcal{K},\quad r\geq 0\end{split}$ (12)
@@ III-C Nesterov-Todd Scaling
    SOC : For $s,\ z\in\text{int}\mathbb{Q}^{d}$ with $d>1$, let $\bar{s}=(\bar{s}_{0},\bar{s}_{1})$, $\bar{z}=(\bar{z}_{0},\bar{z}_{1})\in\mathbb{R}\times\mathbb{R}^{d-1}$ be the normalized points
@@
    and $\gamma$, $\bar{w}=(\bar{w}_{0},\bar{w}_{1})\in\mathbb{R}\times\mathbb{R}^{d-1}$, $\eta$ be
    $\gamma=\left(\frac{1+\bar{s}^{T}\bar{z}}{2}\right)^{1/2}$ (19)
[…]
    $\eta=\left(\frac{s_{0}^{2}-s_{1}^{T}s_{1}}{z_{0}^{2}-z_{1}^{T}z_{1}}\right)^{1/4}$ (21)
@@
$W=\texttt{blkdiag}(W_{1},\cdots,W_{l+n_{SOC}})$ (23)
@@ III-D Search Directions
The solver developed in this paper utilizes PDIPMs, which are known to be among the most efficient frameworks within IPMs [46]. PDIPMs compute a Newton-like search direction by linearizing the (scaled) central path equations and loosely follow the central path toward the solution. Since the intermediate iterates of PDIPMs do not lie exactly on the central path, they are not generally primal-dual feasible to ($\mathcal{P}$) and ($\mathcal{D}$), as well as equations (11b), (11c) or (16) are not precisely satisfied for the barrier parameter $\beta$ [60, Chapter 10]. Therefore, duality gap cannot be evaluated as $s^{T}z+\kappa\tau$ at intermediate iterates. Instead, PDIPMs introduce the duality measure $\mu$ as an indicator of the approximated duality gap at each point in the search space instead of $\beta$ [60, 46].
$\mu=\frac{s^{T}z+\kappa\tau}{l+n_{SOC}+1}>0$ (24)
@@
$(W^{-T}s)\circ(Wz)=\sigma\mu e,\quad\kappa\tau=\sigma\mu e$ (25)
@@
$\Lambda:=\begin{bmatrix}Q&A^{T}&G^{T}&q\\ -A&0&0&b\\ -G&0&0&h\\ -q^{T}-\frac{2}{\tau}x^{T}Q&-b^{T}&-h^{T}&\frac{1}{\tau^{2}}x^{T}Qx\end{bmatrix}$ (27)
@@
$\mu_{a}=\frac{(s+\alpha_{a}\Delta s_{a})^{T}(z+\alpha_{a}\Delta z_{a})+(\kappa+\alpha_{a}\Delta\kappa_{a})(\tau+\alpha_{a}\Delta\tau_{a})}{l+n_{SOC}+1}$ (31)
and $\alpha_{a}$ is the step size obtained from the line search in (32) along the affine search direction (i.e., by replacing $\Delta(s,z,\kappa,\tau)$ in (32) with $\Delta(s_{a},z_{a},\kappa_{a},\tau_{a})$).
$\begin{split}\alpha_{\text{step}}=\sup\{\bar{\alpha}\in(0,1)\left.\right|\ &(s,z)+\bar{\alpha}(\Delta s,\Delta z)\in\text{int}\mathcal{K}\times\text{int}\mathcal{K}^{\ast},\\ &(\kappa,\tau)+\bar{\alpha}(\Delta\kappa,\Delta\tau)>0\}\end{split}$ (32)
@@
where the step size $\alpha$ is obtained from the line search in (32) along the combined search direction $\Delta(s_{c},z_{c},\kappa_{c},\tau_{c})$.
@@
$\Delta\xi$ $:=\begin{bmatrix}\Delta x^{T}&\Delta y^{T}&\Delta z^{T}\end{bmatrix}^{T}$ (37a)
$b_{1}$ $:=\begin{bmatrix}-q^{T}&b^{T}&h^{T}\end{bmatrix}^{T}$ (37b)
$b_{2}$ $:=\begin{bmatrix}-d_{x}^{T}&d_{y}^{T}&(d_{z}-W(\lambda\backslash d_{s}))^{T}\end{bmatrix}^{T}$ (37c)
@@
$K:=\begin{bmatrix}Q&A^{T}&G^{T}\\ A&0&0\\ G&0&-W^{T}W\end{bmatrix}$ (38)
@@
$\Delta\xi=\Delta\xi_{2}+\Delta\tau\Delta\xi_{1}$ (41)
@@
$\Delta s$ $=W(\lambda\backslash d_{s}-W\Delta z)$ (42a)
$\Delta\kappa$ $=(d_{\kappa}-\kappa\Delta\tau)/\tau$ (42b)
@@
To further improve computational efficiency in matrix factorization, we adopt the structure exploitation strategy for SOC scaling proposed in [20] to obtain a sparse KKT matrix. For a SOC $\mathbb{Q}^{d}$, the scaling matrix $W_{SOC}$ in (22b) propagates to the (3, 3) block in the KKT matrix as $W^{T}_{SOC}W_{SOC}$, which is a dense matrix with $d^{2}$ nonzero entries. This results in a significant number of nonzeros in both the KKT matrix and its factorization, even for problems with sparse data. By exploiting the diagonal-plus-rank-one structure of $W_{SOC}$, we can expand $W^{T}_{SOC}W_{SOC}$ to a sparse form as
$\tilde{W}_{SOC}^{T}\tilde{W}_{SOC}=\eta^{2}\begin{bmatrix}\bar{D}&v&u\\ v^{T}&1&0\\ u^{T}&0&-1\end{bmatrix}$ (43)
@@
$\bar{D}:=\begin{bmatrix}a&\\ &I_{d-1}\end{bmatrix},\ u:=\begin{bmatrix}u_{0}\\ u_{1}\bar{w}_{1}\end{bmatrix},\ v:=\begin{bmatrix}0\\ v_{1}\bar{w}_{1}\end{bmatrix}$